Two types of rice varieties are being considered for yield and a comparison is needed. 28 hectares were planted with rice varieties exposed to fairly uniform growing conditions. The results are tabulated below:


Variety A Variety B

Average yield 68 sacks/ha 52 sacks/ha

Sample Variance 7.2 10.4


Is there a significant difference between variety A and variety B @ 5% significant level.


1
Expert's answer
2021-02-23T09:07:53-0500

We should use t-test:

1) H0: "\mu_1=\mu_2" and H1: "\mu_1\ne\mu_2"

2) "\alpha=0.05"; two-tailed; d.f. = 28-1 = 27

critical values = 2.052 and -2.052

3) "t=\frac{(\overline{X}_1-\overline{X}_2)-(\mu_1-\mu_2)}{\sqrt{s^2_1/n_1+s^2_2/n_2}}=\frac{68-52}{\sqrt{7.2/28+10.4/28}}=20.18"

4) 20.18>2.052, so we should reject the null hypothesis

5) There is a significant difference between variety A and variety B